henry paid $.80 for each bag of peanuts. he sold all but 20 of them for $1.50 and made a profit of $54. how many bags did he buy?
Answer by mananth(16949)

cost each bag = $0.8
number of bags he bought = x
he sold x-20 bags
he sold at $1.50
...
he made a profit of $54
..
1.50(x-20)-0.8*x=54
1.50x-30-0.8x=54
0.7x=54+30
/0.7
x=84/0.7
x=120 bags of peanuts
...
m.ananth@hotmail.ca